# HGCLIP: Exploring Vision-Language Models with Graph Representations for Hierarchical Understanding

### v1

@@ 4.3 Delving into Graph Representations
[…]
@@
[…]
@@
5 Experiment
5.1 Benchmark Setting
Hierarchical Image Classification: We consider both generic (CIFAR-100 [26] and Caltech-101 [8]) and fine-grained (FGVC-Aircraft [30], StanfordCars [25], Food-101 [1], Fruits-360 [31], OxfordPets-37 [34] and ETHEC [6]) visual classification tasks to demonstrate our method under general situations of data diversity, where the label hierarchical levels range from two to four.
Domain Generalization: We evaluate the robustness of our approach on out-of-distribution datasets. The source distributions correspond to the original ImageNet [5]. The task is to classify images from the target datasets (ImageNetV2 [37], ImageNet-Sketch [44], ImageNet-A [17] and ImageNet-R [16]), which consist of images that contain various types of domain shifts. It is important to note that due to the inconsistent semantic granularity at each hierarchical level in the original ImageNet, we only select all the categories from ImageNet-A and ImageNet-R (a total of 314 categories) for our experiments, opting only for categories from two adjacent hierarchical levels.
Subpopulation Shift: We also examine HGCLIP’s robustness to subpopulation shift within a dataset. The source and target distributions, though encompassing identical class categories, feature distinct subpopulations within those classes. Our empirical investigations were executed on the four BREEDS [39] ImageNet subsets: Living17, Nonliving26, Entity13, and Entity30.
Implementation Details: We use top-1 accuracy to evaluate the prediction performance. We adopt CLIP ViT-B/16 as the visual encoder and use the corresponding CLIP Transformer as the text encoder, where $D=512$. We set $\lambda_{1}$ = 1 and $\lambda_{2}$ = 0.2 to weight the proportion of hierarchical structural information. For hierarchical classification, we use deep prompting with $v=t=4$ in the first 9 transformer layers and train the models for 50 epochs. For distribution shift benchmarks, we use the same number of prompt tokens in the first 3 layers and train the models for 20 epochs. All models are trained with a batch size of 64 and a learning rate of 3e-4 via SGD optimizer, and decay by the cosine annealing rule during training. All the experiments are implemented on PyTorch 1.13.0 using an NVIDIA GeForce RTX 4090 GPU. For additional benchmark setting and implementation details, see Appendix A.
5.2 Hierarchical Image Classification
Table 1 displays the performance of HGCLIP on 8 hierarchical classification datasets, including 2 generic and 6 fine-grained. We compare its performance with zero-shot CLIP and recent works on prompt learning. In the case of CLIP, we utilize handcrafted specific prompts designed for each dataset. In comparison with state-of-the-art MaPLe [22] and PromptSRC [23], HGCLIP exhibits improved performance across all levels on all the datasets, with the exception of a slight decline in performance on Fruits-360. With the contribution of graph representations, as opposed to MaPLe and PromptSRC, HGCLIP demonstrates superior generalization across multiple hierarchical categories on all the datasets, achieving an absolute average gain of 2.1% and 6.3% respectively.
@@ 5.2 Hierarchical Image Classification
Table 1: Top-1 accuracy (%) comparison on hierarchical image classification of HGCLIP with previous CLIP-based prompt tuning methods. The best result is bold and the second best is underlined. * denotes that the dataset is with available class hierarchy, and hierarchies of others are queried through ChatGPT [33]. $l_{i}$ represents the classification accuracy at the $i$-th hierarchical level, where a smaller $i$ indicates a coarser granularity level, and vice versa.
@@
CLIP-Adapter [10] IJCV’23 83.91 77.03
Tip-Adapter [53] ECCV’22 84.57 81.42
CALIP [14] AAAI’23 77.51 74.28
@@
Source Target
Method ImageNet* -V2* -Sketch* -A -R
CLIP 86.11 80.19 72.13 46.11 70.85
CoOp 85.44 79.71 70.85 45.63 70.91
CoCoOp 85.03 79.49 72.96 46.99 72.47
MaPLe 91.22 85.12 74.19 49.28 73.79
PromptSRC 89.89 83.79 75.04 49.37 74.22
HGCLIP 92.19 86.24 77.40 50.38 76.07
Table 3: Domain generalization. Comparison with CLIP-based methods on robustness to domain shifts. * denotes that 314 categories out of 1,000 categories are selected for the dataset.
Module CIFAR-100 FVGC-Aircraft
@@
TP and VP serve as textual and visual prompts. TG and VG denote graph encoder for textual and visual modality respectively.
Table 4: Component Analysis of HGCLIP.
Moreover, in Table 2, we compare HGCLIP with prior feature adaption methods based on CLIP. CLIP-Adapter [10] learns two residual-style adapters after CLIP. Tip-Adapter [53] constructs a key-value cache model by extracting features from few-shot data through CLIP, then views the cache model as a well-performing initialization and fine-tunes the cache keys. CALIP [14] is proposed to boost CLIP performance via a parameter-free attention module between multi-modal representations. In comparison with these feature adaption approaches, HGCLIP exhibits excellent feature representation capabilities, with an accuracy on CIFAR-100 that is 8.7%, 6.2%, and 13.3% higher than theirs, respectively.
5.3 Distribution Shifts
Domain Generalization: Table 3 summarizes the results of HGCLIP and prior approaches on out-of-distribution datasets. We verify the transferability of models trained on ImageNet to various ImageNet variants with domain shifts. On the target datasets, the performance of HGCLIP surpasses previous SoTA methods. This achievement underscores the efficacy of hierarchical graph representations to enhance multi-modal features. Such an integration improves the generalization capabilities, enabling it to perform well across varying domains. This indicates that HGCLIP not only captures the intricate relationships within the data but also adapts effectively to new and unseen domains.
Subpopulation Shift: We further evaluate the generalizability of HGCLIP on four datasets from BREEDS (subsets of ImageNet) that exhibit subpopulation shifts and provide available class hierarchies. Figure 5 depicts the performance of HGCLIP and previous methods on this subpopulation shift benchmark. The models are trained only on base classes and tested on novel classes. The results suggest that HGCLIP possesses strong generalization capabilities even when confronted with feature-level shifts, underscoring the efficacy of hierarchical structure information in enhancing model generalizability. This success demonstrates the robustness of HGCLIP in handling variations within subpopulations, ensuring consistent accuracy and reliability across diverse and shifting data landscapes.
Figure 5: Results for different methods on the four BREEDS datasets [39] to measure robustness to subpopulation shift.
@@ 5.3 Distribution Shifts
5.4 Ablative Analysis
Components Analysis: HGCLIP primarily consists of multi-modal prompts and graph encoders. In Table 4, we ablate on the performance of each module. The prompts facilitate the model in learning hierarchical contextual features, while the graph encoders effectively integrate hierarchical structure information into the feature representations. This enables the model to achieve impressive results across multiple granularities of hierarchical categories.
Calculating Logits for Hierarchical Classification: In this context, there are two approaches for calculating hierarchical logits. The first approach involves calculating the logits for each category separately based on the hierarchical relationships. The second approach is bottom-up: logits for only the finest-grained categories are initially computed. Logits for the next higher level of categories are then derived by summing up the logits of their corresponding subordinate categories. We compare the performance of these two computation methods in zero-shot experiments and find that while maintaining consistent prompt templates, different hierarchical computation methods yielded varying results at the coarse level, with the former reaching 43.22% and the latter only achieving 6.13% on CIFAR-100. Refer to Appendix B for additional illustration.
Noisy Hierarchies Queried by Large Language Models: It is important to note that LLMs may output sub-optimal hierarchical labels. LLMs produce inconsistent hierarchical labels based on a set of input category names or generate hierarchical labels of different levels, leading to certain biases in the model performance. However, even when utilizing the noisy hierarchy, HGCLIP still enhances accuracy within the original categories in the dataset. For more analysis of sub-optimal hierarchical labels, see Appendix C.
@@ 5.6 Graph Encoder Analysis
We further conduct experiments to analyze the impact of various graph encoders. We apply three of the most commonly used graph learning models: GCN [24], GAT [43], and GraphSAGE [15] to HGCLIP, the results are illustrated in Figure 7. First, for both hierarchical levels, GAT consistently exhibits superior performance, particularly at the fine level, where GAT markedly surpasses the other two graph encoders. Second, with the increase of layer depth of the graph encoder, the accuracy initially rises. Upon reaching a peak (3 layers), the accuracy begins to gradually decline with further increase in layer depth. Therefore, we use a 3-layer GAT as graph encoder in our experiments.
Figure 7: The impact of different GNN variants and the number of layers on the performance of hierarchical classification. We report accuracy results on two hierarchical levels.
@@ 6 Conclusion
In this work, we propose a novel view that combines VLMs with graph representations. Deep prompts are incorporated into the multi-modal branches, allowing VLMs to better learn hierarchical representations. The graph-based hierarchical relationships are encoded into features, strengthening the connection of features across multiple granularities. When integrating image features with graph representation, given that image features are pixel/region-level, prototype learning is employed for class-level image features, which are then fused with the original image features through the attention mechanism. HGCLIP achieves SoTA results on several hierarchical image classification benchmarks and demonstrates robustness and generalization on multiple datasets with distribution shifts.
Supplementary Material
Figure 8: Illustrations of GPT-generated hierarchical labels.
The following section contains supplemental information and encompasses more implementation details, results comparison, and a thorough ablative analysis of HGCLIP. The contents are organized in the following order.
  • Additional Implementation Details (Appendix A)
  • Additional Method Introduction (Appendix B)
  • Additional Ablative Analysis (Appendix C)
  • Additional Results Presentation (Appendix D)
Appendix A Implementation Details
@@
Hierarchical classes number: Table 5 presents the number of categories at each hierarchical level for all the datasets utilized in the experiments.
Fruits-360 [31]: From the perspective of class names, its original form is somewhat incomplete, as some classes are distinguished only by numerical indices (e.g., “Apple Golden 1” and “Apple Golden 2”). Therefore, we manually rename the classes using the structure listed in Table 6, which leads to a reduction in the number of subclasses from 131 to 113. This makes the evaluation of model performance more reasonable and accurate.
BREEDS [39]: Since our method necessitates the use of class hierarchies, we opt for the ’good’ split of subpopulations within a given superclass into source and target domains, signifying that subpopulations in the source and target domains share a common parent. In contrast, the other splits i.e., ’bad’ and ’random’ would result in a greater degree of separation between the subpopulations of the source and target domains in terms of hierarchical structure distance. However, such an arrangement would not effectively validate the efficacy of our method. Although these two ways of partitioning the subpopulations would lead to stronger adversarial conditions, it would cause the model to learn the wrong label hierarchy. Thus, they are not considered in our analysis.
# Import Necessary Libraries
from robustness.tools.breeds_helpers
import BreedsDatasetGenerator
# Creating BREEDS datasets
DG = BreedsDatasetGenerator(info_dir)
# Example: Balanced dataset from level=2 with
# subclasses partitioned in "good"
ret = DG.get_superclasses(level=2,
                      Nsubclasses=None,
                      split="good",
                      ancestor=None,
                      balanced=True)
superclasses, subclass_split, label_map = ret
Textual diversity: Table 7 displays the hand-crafted prompts used for textual prompt initialization.
Implementation Details: We use a publically available ViT-B/16 CLIP model with $D$ = 512 and use a learning rate of 3e-4 which is fixed for all experiments in all benchmarks. We set $\lambda_{1}$ = 1 and $\lambda_{2}$ = 0.2 to weight the proportion of hierarchical structural information. The corresponding hyperparameters are fixed across all datasets and benchmark. The respective epochs are fixed across all datasets. For hierarchical classification, we use deep prompting with $v=t=4$ in the first 9 transformer layers and train for 50 epochs. Since the classification of coarse-grained categories is noticeably less challenging than that of fine-grained ones, we set $w_{h}$ in Eq. 13 to 2 and the remaining $\{w_{i}\}_{i=1}^{h-1}$ to 1, in order to improve the model performance on fine-grained classification. For distribution shift benchmarks, we use the same number of prompt tokens in the first 3 layers and train for 20 epochs. All models are trained using SGD optimizer and utilize a single NVIDIA GeForce RTX 4090 GPU.
@@
Original Class Cleaned Subclass
Apple Golden 1 Apple Golden
Apple Golden 2 Apple Golden
Apple Golden 3 Apple Golden
Apple Red 1 Apple Red
Apple Red 2 Apple Red
Apple Red 3 Apple Red
Apple Red Yellow 1 Apple Red Yellow
Apple Red Yellow 2 Apple Red Yellow
Cantaloupe 1 Cantaloupe
Cantaloupe 2 Cantaloupe
Cherry 1 Cherry
Cherry 2 Cherry
Cucumber Ripe 2 Cucumber Ripe
Grape White 2 Grape White
Grape White 3 Grape White
Grape White 4 Grape White
Peach 2 Peach
Pear 2 Pear
Plum 2 Plum
Plum 3 Plum
Tomato 1 Tomato
Tomato 2 Tomato
Tomato 3 Tomato
Tomato 4 Tomato
Table 6: Mapping from original class names to new class names for Fruits-360.
@@
FVGC-Aircraft a photo of an aircraft [CLASS].
@@
ImageNetV2 [37] a photo of a [CLASS].
ImageNet-Sketch [44] a photo of a [CLASS].
ImageNet-A [17] a photo of a [CLASS].
ImageNet-R [16] a photo of a [CLASS].
@@
Appendix B Method Introduction

### v3

@@ 4.3 Delving into Graph Representations
$\textit{logits}=\lambda_{1}\cdot\tilde{\mathbf{f}_{v}}{\hat{\mathbf{F}_{t}}}^{T}+\lambda_{2}\cdot\hat{\mathbf{f}_{v}}{\hat{\mathbf{F}_{t}}}^{T},$ (12)
@@
For hierarchical image classification, the model is required to simultaneously predict several labels at different granularities. Consequently, it is necessary to partition the predicted logits into their respective hierarchical categories $\textit{logits}_{i}$, with each level corresponding to the ground truth labels $GT_{i}$, where $i=1,\cdots,h$. The overall loss function can be defined as follows
$\mathcal{L}=\sum_{i=1}^{h}w_{i}\cdot\mathcal{L}_{CE}({GT}_{i},\textit{logits}_{i}),$ (13)
@@
Table 1: Top-1 accuracy (%) comparison on hierarchical image classification of HGCLIP with previous CLIP-based prompt tuning methods. The best result is bold and the second best is underlined. * denotes that the dataset is with available class hierarchy, and hierarchies of others are queried through ChatGPT. $l_{i}$ represents the classification accuracy at the $i$-th hierarchical level, where a smaller $i$ indicates a coarser granularity level, and vice versa.
5 Experiment
5.1 Benchmark Setting
Hierarchical Image Classification: We consider 11 visual classification datasets, covering a wide range of recognition tasks. These include two general object datasets, CIFAR-100 Krizhevsky et al. (2009) and Caltech-101 Fei-Fei et al. (2004); six fine-grained datasets, FGVC-Aircraft Maji et al. (2013), StanfordCars Krause et al. (2013), Food-101 Bossard et al. (2014), Fruits-360 Mureşan and Oltean (2017), OxfordPets-37 Parkhi et al. (2012) and ETHEC Dhall et al. (2020); a scene recognition dataset SUN397 Xiao et al. (2010); a texture dataset DTD Cimpoi et al. (2014) and a satellite image dataset EuroSAT Helber et al. (2019). The aim is to demonstrate our method under general situations of data diversity, where the label hierarchical levels range from two to four.
Implementation Details: We use top-1 accuracy to evaluate the prediction performance. We adopt CLIP ViT-B/16 as the visual encoder and use the corresponding CLIP Transformer as the text encoder. We set $\lambda_{1}$ = 1 and $\lambda_{2}$ = 0.2 to weight the proportion of hierarchical structural information. For hierarchical classification, we use deep prompting with $v=t=4$ in the first 9 transformer layers and train the models for 50 epochs. All models are trained with a batch size of 64 and a learning rate of 3e-4 via SGD optimizer, and decay by the cosine annealing rule during training.
5.2 Hierarchical Image Classification
CLIP-based prompt tuning methods. Table 1 displays the comparative performance of zero-shot CLIP, recent works on prompt learning and HGCLIP on 11 diverse hierarchical classification datasets. In the case of CLIP, we utilize handcrafted specific prompts designed for each dataset. In comparison with state-of-the-art MaPLe Khattak et al. (2023a) and PromptSRC Khattak et al. (2023b), HGCLIP exhibits improved performance across all levels on all the datasets, with the exception of a slight decline in performance on Fruits-360. With the contribution of graph representations, as opposed to SoTA MaPLe and PromptSRC, HGCLIP demonstrates superior generalization across multiple hierarchical categories on all the datasets, achieving an absolute average gain of 2.2% and 5.7% respectively.
CLIP-based feature adaptation methods. In Table 2, we compare HGCLIP with prior feature adaption methods based on CLIP. CLIP-Adapter Gao et al. (2023) learns two residual-style adapters after CLIP. Tip-Adapter Zhang et al. (2022) constructs a key-value cache model by extracting features from few-shot data, then views the cache model as a well-performing initialization and fine-tunes the cache keys. CALIP Guo et al. (2023) is proposed to boost CLIP performance via a parameter-free attention module between multi-modal representations. In comparison with these feature adaption approaches, HGCLIP exhibits excellent feature representation capabilities, with an accuracy on CIFAR-100 that is 8.7%, 6.2%, and 13.3% higher than theirs, respectively.
Visual-only hierarchical image classification methods. We have analysed various multi-modal methods above, and to demonstrate the effectiveness of HGCLIP, we compare visual-only fine-grained visual classification methods, as shown in Table 3. Our method still achieve a significant advantage. Additionally, the visual-only FGVC methods are more time-consuming compared to ours (100 v.s. 50 training epochs).
@@ 5.2 Hierarchical Image Classification
CLIP-Adapter Gao et al. (2023) IJCV’23 83.91 77.03
Tip-Adapter Zhang et al. (2022) ECCV’22 84.57 81.42
CALIP Guo et al. (2023) AAAI’23 77.51 74.28
@@
Method Acc. %
$l_{1}$ $l_{2}$
PMG Du et al. (2020) ECCV’20 87.16 83.02
FGN Chang et al. (2021) CVPR’21 87.88 83.60
GHORD Zhao et al. (2021) CVPR’21 87.93 84.36
CHRF Liu et al. (2022) ECCV’22 88.67 84.91
TFGIC Xu et al. (2023) AAAI’23 89.20 85.17
HGCLIP (Ours) 91.87 86.55
Table 3: Comparison with visual-only SOTA FGVC methods.
Module CIFAR-100 FGVC-Aircraft
@@
Table 4: Component Analysis of HGCLIP. TP and VP serve as textual and visual prompts. TG and VG denote graph encoder for textual and visual modality.
5.3 Ablative Analysis
Components Analysis. HGCLIP primarily consists of multi-modal prompts and graph encoders. In Table 4, we ablate on the performance of each module. The prompts facilitate the model in learning hierarchical contextual features, while the graph encoders effectively integrate hierarchical structure information into the feature representations. This enables the model to achieve impressive results across multiple semantic granularities.
Noisy Hierarchies Queried by LLMs. It is important to note that LLMs may output sub-optimal hierarchical labels. LLMs produce inconsistent hierarchical labels based on a set of input category names or generate hierarchical labels of different levels, leading to certain biases in the model performance. However, even when utilizing the noisy hierarchy, HGCLIP still enhances accuracy within the original categories in the dataset.
@@ 5.5 Graph Encoder Analysis
We further conduct experiments to analyze the impact of various graph encoders. We apply three of the most commonly used graph learning models: GCN Kipf and Welling (2016), GAT Veličković et al. (2018b), and GraphSAGE Hamilton et al. (2017) to HGCLIP, the results are illustrated in Figure 6. First, for both hierarchical levels, GAT consistently exhibits superior performance, particularly at the fine-grained level, where GAT surpasses the other encoders. Second, with the increase of layer depth of the graph encoder, the accuracy initially rises. Upon reaching a peak (3 layers), the accuracy begins to gradually decline with further increase in layer depth. Therefore, we use a 3-layer GAT as graph encoder in our experiments.
Figure 6: The impact of various GNN variants and the number of layers on hierarchical classification. We report accuracy results on two hierarchical levels.
@@ 6 Conclusion
In this work, we propose a novel view that combines VLMs with graph representations. Deep prompts are incorporated into the multi-modal branches, allowing VLMs to better learn hierarchical representations. The graph-based hierarchical relationships are encoded into features, strengthening the connection of features across multiple granularities. When integrating image features with graph representation, given that image features are pixel/region-level, prototype learning is employed for class-level image features, which are then fused with the image features through the attention mechanism. HGCLIP achieves SoTA results on several hierarchical image classification benchmarks.
Limitations
Although this work has achieved results by utilizing graph representations to characterize hierarchical information, it merely employs a simple graph encoder (such as GCN, GAT, GraphSAGE) to encode structural information. With the advancement of graph learning, it is anticipated that there will be better graph learning methods for representing hierarchical structures, which could further enhance features. We separately employ graph encoders for different modalities, yielding the best performance, outperforming cross-modal graph encoders with shared weights. This may be because current multi-modal models lack the ability to accurately extract features from each modality. With the development of multimodal models, this limitation is expected to be addressed, significantly reducing training time and improving inference speed. Additionally, we leverage multi-modal prompt learning but have not integrated it with graph learning, perhaps combining the two could yield better results.
Appendix A Data Details
A.1 Hierarchical Classes Number
@@
Table 5 presents the number of categories at each hierarchical level for all the datasets utilized in the experiments.
@@
A.2 Textual Diversity
Table 6 displays the hand-crafted prompts used for textual prompt initialization.
@@ A.2 Textual Diversity
FGVC-Aircraft a photo of an aircraft [CLASS].
@@
SUN397 a photo of [CLASS].
DTD a photo of texture [CLASS].
EuroSAT a centered satellite photo of [CLASS].
@@
ImageNetV2 a photo of a [CLASS].
ImageNet-Sketch a photo of a [CLASS].
ImageNet-A a photo of a [CLASS].
ImageNet-R a photo of a [CLASS].
@@
A.3 Dataset Setting
Fruits-360 Mureşan and Oltean (2017): From the perspective of class names, its original form is somewhat incomplete, as some classes are distinguished only by numerical indices (e.g., “Apple Golden 1" and “Apple Golden 2"). Therefore, we manually rename the classes, which leads to a reduction in the number of subclasses from 131 to 113. This makes the evaluation of model performance more reasonable and accurate.
BREEDS Santurkar et al. (2021): Since our method necessitates the use of class hierarchies, we opt for the ’good’ split of subpopulations within a given superclass into source and target domains, signifying that subpopulations in the source and target domains share a common parent. In contrast, the other splits i.e., ’bad’ and ’random’ would result in a greater degree of separation between the subpopulations of the source and target domains in terms of hierarchical structure distance. However, such an arrangement would not effectively validate the efficacy of our method. Although these two ways of partitioning the subpopulations would lead to stronger adversarial conditions, it would cause the model to learn the wrong label hierarchy. Thus, they are not considered in our analysis.
Appendix B Implementation Details
We use a publically available ViT-B/16 CLIP model with $D$ = 512 and use a learning rate of 3e-4 which is fixed for all experiments in all benchmarks. We set $\lambda_{1}$ = 1 and $\lambda_{2}$ = 0.2 to weight the proportion of hierarchical structural information. The corresponding hyperparameters are fixed across all datasets and benchmark. The respective epochs are fixed across all datasets. For hierarchical classification, we use deep prompting with $v=t=4$ in the first 9 transformer layers and train for 50 epochs. Since the classification of coarse-grained categories is noticeably less challenging than that of fine-grained ones, we set $w_{h}$ in Eq. 13 to 2 and the remaining $\{w_{i}\}_{i=1}^{h-1}$ to 1, in order to improve the model performance on fine-grained classification. For distribution shift benchmarks, we use the same number of prompt tokens in the first 3 layers and train for 20 epochs. All models are trained using SGD optimizer and utilize 4 NVIDIA GeForce RTX 4090 GPUs.
Appendix C Distribution Shifts
Domain Generalization: We evaluate the robustness of our approach on out-of-distribution datasets Xia et al. (2024b); Hu et al. (2024). The source distributions correspond to the original ImageNet Deng et al. (2009). The task is to classify images from the target datasets (ImageNetV2 Recht et al. (2019), ImageNet-Sketch Wang et al. (2019), ImageNet-A Hendrycks et al. (2021b) and ImageNet-R Hendrycks et al. (2021a)), which consist of images that contain various types of domain shifts. It is important to note that due to the inconsistent semantic granularity at each hierarchical level in the original ImageNet, we only select all the categories from ImageNet-A and ImageNet-R to facilitate our experiment. Table 7 summarizes the results of HGCLIP and prior approaches on out-of-distribution datasets. We verify the transferability of models trained on ImageNet to various ImageNet variants with domain shifts. On the target datasets, the performance of HGCLIP surpasses previous SoTA methods. This achievement underscores the efficacy of hierarchical graph representations to enhance multi-modal features. Such an integration improves the generalization capabilities, enabling it to perform well across varying domains. This indicates that HGCLIP not only captures the intricate relationships within the data but also adapts effectively to new and unseen domains.
Subpopulation Shift: We also examine HGCLIP’s robustness to subpopulation shift within a dataset. The source and target distributions, though encompassing identical class categories, feature distinct subpopulations within those classes. Our empirical investigations were executed on the four BREEDS Santurkar et al. (2021) ImageNet subsets: Living17, Nonliving26, Entity13, and Entity30. We further evaluate the generalizability of HGCLIP on four datasets from BREEDS (subsets of ImageNet) that exhibit subpopulation shifts and provide available class hierarchies. Figure 7 depicts the performance of HGCLIP and previous methods under subpopulation shifts. The models are trained only on base classes and tested on novel classes. The results suggest that HGCLIP possesses strong generalization capabilities even when confronted with feature-level shifts, underscoring the efficacy of hierarchical structure information in enhancing model generalizability. This success demonstrates the robustness of HGCLIP in handling variations within subpopulations, ensuring consistent accuracy and reliability across diverse and shifting data landscapes.
Figure 7: Results for different methods on the four BREEDS datasets Santurkar et al. (2021) to measure robustness to subpopulation shift.
Source Target
Method ImageNet -V2 -Sketch -A -R
CLIP 86.11 80.19 72.13 46.11 70.85
CoOp 85.44 79.71 70.85 45.63 70.91
CoCoOp 85.03 79.49 72.96 46.99 72.47
MaPLe 91.22 85.12 74.19 49.28 73.79
PromptSRC 89.89 83.79 75.04 49.37 74.22
HGCLIP 92.19 86.24 77.40 50.38 76.07
Table 7: Domain generalization. Comparison with CLIP-based methods on robustness to domain shifts.
Appendix D Method Introduction
D.1 Calculating Logits for Hierarchical Classification